# Gravitational wave polarization modes and the kinematical tensors in general relativity and beyond

### v1

@@ Abstract
Interrelationships between the expansion, the shear, and the vorticity—the kinematical tensors—on the one hand, and the polarization modes of gravitational waves on the other hand, are studied by considering freely falling test particles. After studying exact relations, we consider slowly moving particles under the influence of a weak gravitational field. Linearized plane waves of metric theories of gravity representative of those determined by a general second order Lagrangian, including General Relativity, are shown to exemplify the following interconnections: between the transverse components of the shear and the transverse tensor polarization mode; between the expansion, and both the transverse scalar and the longitudinal polarization modes; and between the longitudinal-transverse components of the shear and the vorticity, and the vector polarization mode.
Gravitational waves, Polarization modes, Kinematical tensors, Metric theories of gravity
@@ I Introduction
written here in terms of the tidal tensor $K$. Then, measurements of the relative velocity between the particles reveals information about the kinematical tensors, namely $\theta$, $\sigma$, and $\omega$, from which some aspects about the gravitational field could be inferred. Similarly, measuring the relative acceleration reveals information about the part of the Riemann tensor encoded in the tidal tensor—its so-called electric part.
@@
For a general gravitational field, the six independent components of the tidal tensor have different effects on a cloud of test particles. This is particularly relevant to detect a gravitational wave. Given a wave with definite propagation direction, the tidal tensor is uniquely decomposed into its purely longitudinal, mixed longitudinal-transverse, and purely transverse parts, the latter being further decomposed into its tensor and scalar parts. These constitute the polarization modes of the wave [3, 1]. GR predicts that a vacuum propagating wave has only the two polarization modes of the transverse tensor part, but many other metric theories of gravity allow more polarization modes (cf. §11.4.2 of [1], and [4]).
@@
Likewise, the study of the relative velocity of neighboring particles has its own relevance for some types of gravitational fields, standing out cosmological models for which it has been a key element in their construction (cf. §5.1 of [5, *ellis2009republication] and §9.1 of [7]). However, it is not so common in the description of gravitational waves, with notable exceptions [8, 9, 10]. Intuition leads one to expect that the kinematical tensors of a cloud of test particles subject to the effects of a gravitational wave are interrelated with the polarization modes of the wave. Concretely, one expects (at least) the following interrelationships: between the transverse components of the shear and the transverse tensor polarization modes; between the longitudinal component of the shear and the longitudinal polarization mode; and between the expansion and the transverse scalar polarization mode. In the present work we analyze this interrelationship for gravitational waves with a definite propagation direction, within the context of GR and other metric theories of gravity, both in the strong and in the weak field regime.
@@
This paper is organized as follows. In §II we review the definitions and the physical meaning of the kinematical tensors, the tidal tensor, and the polarization modes for a cloud of freely falling test particles. Additionally, we write the polarization modes in terms of the kinematical tensors on the one hand, and in terms of the irreducible parts of the Riemann tensor on the other hand. Then, in §III we specialize the analysis to slowly moving particles in a weak gravitational field, obtaining expressions for both the kinematical tensors and the polarization modes at the considered order. Next, in §IV we evaluate the obtained expressions on linearized gravitational plane waves of a selection of metric theories representative of those determined by a Lagrangian that is a general function of the Riemann tensor, and compare them with GR. In appendix A we write exact expressions for the polarization modes in terms of the Newman-Penrose curvature scalars.
@@ 1 Cloud of freely falling test particles
Let us consider a cloud of freely falling—subject only to the influence of gravity—collisionless test particles, [^1] One might think of the cloud representing the test masses of a laser interferometer experiment, the fluid elements of a gas, or the atoms of a resonant bar. hence moving along geodesics. Then their 4-velocity field $u$ satisfies
@@
respectively, where
@@
is the projector orthogonal to $u$. In terms of them, one has
@@
Then, it follows that $D\xi^{a}/\mathrm{d}\tau=\xi^{b}\nabla_{b}u^{a}$, which together with (6) leads to the expression (1) for the relative velocity. One sees that: the expansion $\theta$ amounts to an isotropic rate of change of the relative separation; the shear $\sigma$ corresponds to a volume preserving deformation; and the vorticity $\omega$ amounts to a rotation of the cloud (cf., e.g., §2.5 of [5, *ellis2009republication]).
Moreover, it follow that $\xi$ satisfies the equation of geodesic deviation
@@
leads to the expression (2) for the relative acceleration. The algebraic symmetries of the Riemann tensor imply that $K_{ab}=K_{(ab)}$ and $K_{ab}u^{b}=0$, so $K$ has only six independent components. Furthermore, if one combines (6) with the Ricci identities, an expression for the tidal tensor in terms of the kinematical tensors and their derivatives is obtained. That expression can be simplified using the equations in (3) together with their differential consequences, yielding (cf. §2.3 of [11, *ehlers1993contributions])
$K_{ab}=\frac{1}{3}\left(\frac{\mathrm{d}\theta}{\mathrm{d}\tau}+\frac{\theta^{2}}{3}\right)P_{ab}+\frac{2\theta}{3}\sigma_{ab}+\frac{D\sigma_{ab}}{\mathrm{d}\tau}+\sigma_{ac}{\sigma_{b}}^{c}-\omega_{ac}{\omega_{b}}^{c}.$ (10)
@@ 2 Polarization modes
Now, consider a gravitational wave to which one can associate a definite spacetime propagation direction, for example vacuum GR plane waves, either linearized or exact (cf. §35.3 and §35.9 of [13]). Then, its spatial propagation direction, as seen in the instantaneous rest frame of the test particles, is given by a unit vector $n$ orthogonal to $u$, namely
@@
With it, and with the aid of its orthogonal projector
@@
${K^{\mathrm{l}}}\coloneqq K_{ab}n^{a}n^{b},$ ${K^{\mathrm{s}}}\coloneqq p^{ab}K_{ab},$ ${K^{\mathrm{v}}}_{a}\coloneqq K_{bc}{p_{a}}^{b}n^{c},$ ${K^{\mathrm{tt}}}_{ab}\coloneqq\left({p_{a}}^{c}{p_{b}}^{d}-\frac{1}{2}p_{ab}p^{cd}\right)K_{cd},$ (13)
which we shall call respectively, the longitudinal mode, the transverse scalar mode [^2] Also called breathing mode., the vector mode, and the transverse tensor mode. From their definition, the vector and the transverse tensor modes satisfy ${K^{\mathrm{v}}}_{a}u^{a}={K^{\mathrm{v}}}_{a}n^{a}=0$, ${K^{\mathrm{tt}}}_{ab}u^{b}={K^{\mathrm{tt}}}_{ab}n^{b}=0$, and $g^{ab}{K^{\mathrm{tt}}}_{ab}=0$; thus they have two independent components each. The longitudinal and transverse scalar modes comprise one polarization mode each; while the vector mode and transverse tensor mode comprise two polarization modes each, one for each independent component. In terms of the polarization modes, the tidal tensor gives
@@
The physical effects of the polarization modes stem from the foregoing equation and from (2). A small cloud of test particles initially at rest, tends to: elongate longitudinally due to $K^{\mathrm{l}}$}; expand or compress transversally due to a positive or negative $K^{\mathrm{s}}$}, respectively; elongate along a longitudinal-transverse direction due to $K^{\mathrm{v}}$}; and elongate along a transverse direction due to $K^{\mathrm{tt}}$}. The modes defined here correspond with those defined in the seminal work [3] (cf. their §V A) for linearized plane waves with null propagation as follows: $K^{\mathrm{l}}$} corresponds to their $p_{1}$; the two components ${K^{\mathrm{v}}}_{I}$ correspond to their $p_{2}$ and $p_{3}$; two independent components ${K^{\mathrm{tt}}}_{IJ}$ correspond to their $p_{4}$ and $p_{5}$; and $K^{\mathrm{s}}$} corresponds to their $p_{6}$.
Not all the parts of the Riemann tensor contribute equally to each of the polarization modes. In fact, the irreducible decomposition of the Riemann tensor (cf., e.g., §3.5 of [14]) allows one to write
${K^{\mathrm{v}}}_{a}=\frac{1}{2}S_{bc}{p_{a}}^{b}n^{c}-C_{bcde}{p_{a}}^{b}u^{c}n^{d}u^{e},$ (15c)
${K^{\mathrm{tt}}}_{ab}=\frac{1}{2}S_{cd}{p_{a}}^{c}{p_{b}}^{d}+\frac{1}{4}\left(S_{cd}n^{c}n^{d}-S_{cd}u^{c}u^{d}\right)p_{ab}-C_{cdef}{p_{a}}^{c}u^{d}{p_{b}}^{e}u^{f}-\frac{1}{2}C_{cdef}u^{c}n^{d}u^{e}n^{f}p_{ab},$ (15d)
where $C_{abcd}$ is the Weyl tensor, $S_{ab}=R_{ab}-(R/4)g_{ab}$ is the traceless Ricci tensor, and $R$ is the scalar curvature. It stands out that the scalar curvature does not contribute to the vector and transverse tensor modes.
Similarly, not all kinematical tensors contribute equally to the polarization modes. From the expression for the tidal tensor in terms of the kinematical tensors (10), it follows that
${K^{\mathrm{l}}}=\frac{1}{3}\left(\frac{\mathrm{d}\theta}{\mathrm{d}\tau}+\frac{\theta^{2}}{3}\right)+n^{a}n^{b}\left(\frac{2\theta}{3}\sigma_{ab}+\frac{D\sigma_{ab}}{\mathrm{d}\tau}+\sigma_{ac}{\sigma_{b}}^{c}-\omega_{ac}{\omega_{b}}^{c}\right),$ (16a)
${K^{\mathrm{s}}}=\frac{2}{3}\left(\frac{\mathrm{d}\theta}{\mathrm{d}\tau}+\frac{\theta^{2}}{3}\right)+p^{ab}\left(\frac{2\theta}{3}\sigma_{ab}+\frac{D\sigma_{ab}}{\mathrm{d}\tau}+\sigma_{ac}{\sigma_{b}}^{c}-\omega_{ac}{\omega_{b}}^{c}\right),$ (16b)
@@
${K^{\mathrm{tt}}}_{ab}=\left({p_{a}}^{c}{p_{b}}^{d}-\frac{1}{2}p_{ab}p^{cd}\right)\left(\frac{2\theta}{3}\sigma_{cd}+\frac{D\sigma_{cd}}{\mathrm{d}\tau}+\sigma_{ce}{\sigma_{d}}^{e}-\omega_{ce}{\omega_{d}}^{e}\right).$ (16d)
To gain some insight on the foregoing expressions, it is instructive to analyze the contributions of each kinematical tensor when the others vanish. For example, for a cloud with vanishing shear and vorticity, only the longitudinal and transverse scalar modes survive, determined completely by the combination $\mathrm{d}\theta/\mathrm{d}\tau+\theta^{2}/3$; this reveals a contribution from the expansion to the transverse scalar mode, as expected, but also to the longitudinal mode. On the other hand, for a cloud with vanishing expansion and vorticity, all the modes are nonvanishing and are determined by the following combination of the shear $D\sigma_{ab}/\mathrm{d}\tau+\sigma_{ac}{\sigma_{b}}^{c}$; this shows a contribution from some of its transverse components to the transverse tensor mode and a contribution from its longitudinal component to the longitudinal mode, as anticipated, but also shows a contribution from some of its transverse components to the scalar mode and from its longitudinal-transverse components to the vector mode. Lastly, a rigid cloud, namely one with vanishing expansion and shear, exhibits a similar behavior to that of the preceding example, but with all the modes determined by $\omega_{ac}{\omega_{b}}^{c}$.
@@ 1 Approximation order
thus keeping terms up to order $O(|v^{i}|^{2})$ and $O(h_{ab})$.
@@
where $\dot{h}_{ab}=\partial_{t}h_{ab}$. Incidentally, the foregoing expression reduces to that for the Newtonian limit if the gravitational field is additionally pseudo-stationary, namely if $\dot{h}_{ab}$ is negligible.

### v2

@@ Abstract
Relations between the kinematical tensors (the expansion, the shear, and the vorticity) and the polarization modes of gravitational waves are studied within the context of metric theories of gravity by considering freely falling test particles. After analyzing exact relations, we consider slowly moving particles under the influence of a weak gravitational field. Linearized plane waves of theories representative of those determined by a general second-order Lagrangian, including General Relativity, are shown to exemplify the following interconnections: between the transverse components of the shear and the transverse tensor polarization mode; between the expansion, and both the transverse scalar and the longitudinal polarization modes; and between the longitudinal-transverse components of both the shear and the vorticity, and the vector polarization mode. These relations, from the theoretical point of view, offer a novel insight into the interpretation of the polarization modes in terms of the kinematical tensors. Additionally, the inclusion of the kinematical tensors in the study of gravitational waves might provide a new outlook on their phenomenology.
@@ I Introduction
written here in terms of the tidal tensor $K$. Then, measurements of the relative velocity between the particles reveal information about the kinematical tensors, namely $\theta$, $\sigma$, and $\omega$, from which some aspects about the gravitational field could be inferred. Similarly, measuring the relative acceleration reveals information about the part of the Riemann tensor encoded in the tidal tensor—its so-called electric part.
@@
For a general gravitational field, the six independent components of the tidal tensor have different effects on a cloud of test particles. This is particularly relevant to detect a gravitational wave. Given a wave with definite propagation direction, the tidal tensor is uniquely decomposed into its purely longitudinal, mixed longitudinal-transverse, and purely transverse parts, the latter being further decomposed into its tensor and scalar parts. These constitute the polarization modes of the wave [3, 1]. GR predicts that a vacuum propagating wave has only the two polarization modes of the transverse tensor part, but many other metric theories of gravity allow more polarization modes (cf. §11.4.2 of [1], and [4, 5, 6, 7, 8]). The search for observational evidence of polarization modes beyond GR, is actively pursued by different collaborations (cf. [9, 10, 11, 12] together with §4.3.7 of [13] and references cited therein).
Likewise, the study of the kinematical tensors has its own relevance for some types of gravitational fields, standing out cosmological models for which it has been a key element in their construction (cf. §5.1 of [14, 15] and §9.1 of [16]). However, it is not so common in the description of gravitational waves, with notable exceptions such as [17, 18] and to some extent [19, 20]. Then, the use of a complementary description of the effects of gravitational waves on test particles, such as the one provided by the kinematical tensors, remains largely unexplored and might prove useful.
@@
Following this line of reasoning, in this work we propose studying the kinematical tensors in the context of gravitational waves of metric theories of gravity, how they relate to the polarization modes, and use them to compare some generic modified theories of gravity with GR. Intuition leads one to expect that the kinematical tensors of a cloud of test particles subject to the effects of a gravitational wave are related to the polarization modes of the wave. Concretely, one expects, at least, the following relations: between the transverse components of the shear and the transverse tensor polarization modes; between the longitudinal component of the shear and the longitudinal polarization mode; and between the expansion and the transverse scalar polarization mode. Specifically, we analyze these relations for gravitational waves with a definite propagation direction, within the context of GR and other metric theories of gravity, in both the strong and weak field regimes. To our knowledge, the full set of kinematical tensors had not been used to study gravitational waves.
@@
This paper is organized as follows. In §II we review the definitions and the physical meaning of the kinematical tensors, the tidal tensor, and the polarization modes for a cloud of freely falling test particles. Additionally, we express the polarization modes in terms of the kinematical tensors, and discuss what the kinematical tensors offer to the analysis of gravitational waves. Then, in §III we specialize the analysis to slowly moving particles in a weak gravitational field, obtaining expressions for both the kinematical tensors and the polarization modes at the considered order. Next, in §IV we evaluate the obtained expressions on linearized gravitational plane waves of a selection of metric theories representative of those determined by a Lagrangian that is a general function of the Riemann tensor, and compare them with GR. In appendix A we write exact expressions for the polarization modes in terms the irreducible parts of the Riemann tensor, both in terms of real-valued components and in terms of the Newman-Penrose curvature scalars.
@@ 1 Cloud of freely falling test particles
Let us consider a cloud of freely falling—subject only to the influence of gravity—collisionless test particles, [^1] One might think of the cloud representing the test masses of a laser interferometer experiment, the fluid elements of a gas, or the atoms of a resonant bar. hence moving along geodesics. Then, their 4-velocity field $u$ satisfies
@@
respectively, [^2] In the literature, the vorticity is sometimes defined as the negative of the one defined here. where
@@
is the projector orthogonal to $u$. The kinematical tensors constitute the irreducible decomposition of $\nabla_{a}u_{b}$, namely
@@
Thus, it follows that $D\xi^{a}/\mathrm{d}\tau=\xi^{b}\nabla_{b}u^{a}$, which together with (6) leads to the expression (1) for the relative velocity. One then sees that: the expansion $\theta$ amounts to an isotropic rate of change of the relative separation, namely a change in volume; the shear $\sigma$ corresponds to a volume-preserving deformation; and the vorticity $\omega$ amounts to a rotation of the cloud (cf., e.g., §2.5 of [14, 15]).
Moreover, it follows that $\xi$ satisfies the equation of geodesic deviation
@@
leads to the expression (2) for the relative acceleration between neighboring test particles. The algebraic symmetries of the Riemann tensor imply that $K_{ab}=K_{(ab)}$ and $K_{ab}u^{b}=0$, so $K$ has only six independent components. Furthermore, if one combines (6) with the Ricci identities, an expression for the tidal tensor in terms of the kinematical tensors and their derivatives is obtained; that expression can be simplified using the equations in (3) together with their first covariant derivatives, yielding (cf. §2.3 of [21, 22])
$K_{ab}=\frac{1}{3}\left(\frac{\mathrm{d}\theta}{\mathrm{d}\tau}+\frac{\theta^{2}}{3}\right)P_{ab}+\frac{2\theta}{3}\sigma_{ab}+\frac{D\sigma_{ab}}{\mathrm{d}\tau}\\ +\sigma_{ac}{\sigma_{b}}^{c}-\omega_{ac}{\omega_{b}}^{c}.$ (10)
@@ 2 Polarization modes
Now, we decompose the tidal tensor to obtain the polarization modes of gravitational waves. To do so, consider a gravitational wave to which one can associate a definite spacetime propagation direction, for example vacuum GR plane waves, either linearized or exact (cf. §35.3 and §35.9 of [23]). Then, its spatial propagation direction, as seen in the instantaneous rest frame of the test particles, is given by a unit vector $n$ orthogonal to $u$, namely [^3] An arbitrary observer with 4-velocity $u^{\prime}$ would see a spatial propagation direction $n^{\prime}$ determined by restrictions analogous to those in (11).
@@
Using $n$ and its orthogonal projector
@@
${K^{\mathrm{l}}}\coloneqq K_{ab}n^{a}n^{b},$ (13a)
${K^{\mathrm{s}}}\coloneqq p^{ab}K_{ab},$ (13b)
$K^{\mathrm{v}}_{a}\coloneqq{p_{a}}^{b}K_{bc}n^{c},$ (13c)
$K^{\mathrm{tt}}_{ab}\coloneqq\left({p_{a}}^{c}{p_{b}}^{d}-\frac{1}{2}p_{ab}p^{cd}\right)K_{cd},$ (13d)
which we shall call respectively, the longitudinal mode, the transverse scalar mode [^4] Also called breathing mode., the vector mode, and the transverse tensor mode. From their definition, the vector and the transverse tensor modes satisfy $K^{\mathrm{v}}_{a}u^{a}=K^{\mathrm{v}}_{a}n^{a}=0$, $K^{\mathrm{tt}}_{ab}u^{b}=K^{\mathrm{tt}}_{ab}n^{b}=0$, and $g^{ab}K^{\mathrm{tt}}_{ab}=0$; thus, each of them has two independent components. The longitudinal and transverse scalar modes comprise one polarization mode each; while the vector mode and transverse tensor mode comprise two polarization modes apiece, one for each independent component. In terms of the polarization modes, the tidal tensor gives
@@
The physical effects of the polarization modes stem from the foregoing equation and from the equation of geodesic deviation (2). A small cloud of test particles initially at rest tends to: elongate longitudinally due to $K^{\mathrm{l}}$}; expand or compress transversally due to a positive or negative $K^{\mathrm{s}}$}, respectively; elongate and compress respectively along two mutually orthogonal longitudinal-transverse directions due to $K^{\mathrm{v}}$; and elongate and compress respectively along two mutually orthogonal transverse directions due to $K^{\mathrm{tt}}$. The foregoing can be readily seen if one uses an orthonormal frame adapted to $u$ and $n$ (cf. [2]) or specific coordinate systems adapted to the wave (cf., e.g., §11.4.3 of [1] or [24, 25]). The modes defined here correspond with those defined in the seminal work [3] (cf. their §V A) for linearized plane waves with null propagation as follows: $K^{\mathrm{l}}$} corresponds to their $p_{1}$; the two components $K^{\mathrm{v}}_{I}$ correspond to their $p_{2}$ and $p_{3}$; two independent components $K^{\mathrm{tt}}_{IJ}$ correspond to their $p_{4}$ and $p_{5}$; and $K^{\mathrm{s}}$} corresponds to their $p_{6}$.
Not all the kinematical tensors contribute equally to the polarization modes. From the expression for the tidal tensor in terms of the kinematical tensors (10), it follows that
${K^{\mathrm{l}}}$ $=\frac{1}{3}\left(\frac{\mathrm{d}\theta}{\mathrm{d}\tau}+\frac{\theta^{2}}{3}\right)$
$\quad+n^{a}n^{b}\left(\frac{2\theta}{3}\sigma_{ab}+\frac{D\sigma_{ab}}{\mathrm{d}\tau}+\sigma_{ac}{\sigma_{b}}^{c}-\omega_{ac}{\omega_{b}}^{c}\right),$ (15a)
${K^{\mathrm{s}}}$ $=\frac{2}{3}\left(\frac{\mathrm{d}\theta}{\mathrm{d}\tau}+\frac{\theta^{2}}{3}\right)$
$\quad+p^{ab}\left(\frac{2\theta}{3}\sigma_{ab}+\frac{D\sigma_{ab}}{\mathrm{d}\tau}+\sigma_{ac}{\sigma_{b}}^{c}-\omega_{ac}{\omega_{b}}^{c}\right),$ (15b)
@@
$K^{\mathrm{tt}}_{ab}$ $=\left({p_{a}}^{c}{p_{b}}^{d}-\frac{1}{2}p_{ab}p^{cd}\right)$
$\quad\times\left(\frac{2\theta}{3}\sigma_{cd}+\frac{D\sigma_{cd}}{\mathrm{d}\tau}+\sigma_{ce}{\sigma_{d}}^{e}-\omega_{ce}{\omega_{d}}^{e}\right).$ (15d)
To gain some insight on the foregoing expressions, it is instructive to analyze the contributions of each kinematical tensor when the others vanish. For example, for a cloud with vanishing shear and vorticity, only the longitudinal and transverse scalar modes survive, determined completely by the combination $\mathrm{d}\theta/\mathrm{d}\tau+\theta^{2}/3$; this reveals a contribution from the expansion to the transverse scalar mode, as expected, but also to the longitudinal mode. On the other hand, for a cloud with vanishing expansion and vorticity, all the modes are nonvanishing and are determined by the following combination of the shear $D\sigma_{ab}/\mathrm{d}\tau+\sigma_{ac}{\sigma_{b}}^{c}$; this shows a contribution from some of its transverse components to the transverse tensor mode and a contribution from its longitudinal component to the longitudinal mode, as anticipated, but also shows a contribution from some of its transverse components to the transverse scalar mode and a contribution from its longitudinal-transverse components to the vector mode. Lastly, a rigid cloud, namely one with vanishing expansion and shear, exhibits a similar behavior to that of the preceding example, with all the modes nonvanishing and determined by $\omega_{ac}{\omega_{b}}^{c}$.
Certainly both the kinematical tensors and the polarization modes describe the same physical phenomena: the relative motion of neighboring test particles. In fact, there are some similitudes in the information each of them conveys, but also some dissimilitudes. For example, the volume-preserving deformation due to the shear $\sigma$ resembles the elongation-compression effect resulting from the transverse-tensor mode $K^{\mathrm{tt}}$, while the volume change due to the expansion $\theta$ is analogous to the area change resulting from the transverse scalar mode $K^{\mathrm{s}}$}. However, a rotation of the cloud described by the vorticity $\omega$ cannot be described by a single polarization mode, as discussed in the preceding paragraph. Evidently, the description of the relative motion of test particles can be more transparent either from the kinematical tensors or from the polarization modes depending on the specific type of motion. Thus, the kinematical tensors can be seen as a complementary way of analyzing the physical effects of a gravitational wave. Additionally, they provide a new interpretation of the polarization modes in terms of their effect on the kinematical tensors, by means of the expressions in (15).
@@ 1 Approximation order
thus keeping terms up to order $O(|v^{i}|^{2})$ and $O(h_{ab})$. This is simply a first approximation to the problem. Physical situations where this approximation order applies are those for which the initial velocities of the test particles before the passage of the wave are small compared with the speed of light.
@@
where $\dot{h}_{ab}=\partial_{t}h_{ab}$. Incidentally, the foregoing expression reduces to that for the Newtonian limit if the gravitational field is additionally quasi-stationary, namely if $\dot{h}_{ab}$ is negligible. For subsequent calculations, it is worth recalling the elementary chain rule for derivatives with respect to coordinate time, for example the coordinate acceleration yields $\mathrm{d}v^{i}/\mathrm{d}t=\dot{v}^{i}+v^{k}\partial_{k}v^{i}$ [cf. (18)].